OpenStudy (anonymous):

Factor the polynomial: 54k^2-24

OpenStudy (anonymous):

@jim_thompson5910 Help?

jimthompson5910 (jim_thompson5910):

hint: 54k^2 - 24 6(9k^2 - 4)

OpenStudy (anonymous):

Where did you get the 9k^2 in the second line?

jimthompson5910 (jim_thompson5910):

I factored out 6 from the expression

jimthompson5910 (jim_thompson5910):

54k^2 - 24 6*9k^2 - 6*4 6(9k^2 - 4) .. using the distributive property

OpenStudy (anonymous):

@jim_thompson5910 oh okay I get it, but then there's no number that multipies to give -4 and adds to 9? thats where im stuck at

jimthompson5910 (jim_thompson5910):

from here, you would use the difference of squares rule to factor

jimthompson5910 (jim_thompson5910):

and 9k^2 - 4 is really 9k^2 + 0k - 4 so you need to find two numbers that multiply to 9*(-4) = -36 and add to 0, then you factor by grouping or you can use the difference of squares rule to factor

OpenStudy (anonymous):

the difference of squares rule?

jimthompson5910 (jim_thompson5910):

difference of squares a^2 - b^2 = (a - b)(a+b)

jimthompson5910 (jim_thompson5910):

ex: 25x^2 - 49 (5x)^2 - (7)^2 (5x - 7)(5x + 7) So 25x^2 - 49 factors to (5x - 7)(5x + 7)

OpenStudy (anonymous):